OpenStudy (anonymous):

Solve. sin2x- cosx=1. 0≤x<2π a. 0, 5.07 b. 3.14, 4.32 c. 3.14, 4.36 d. 0.42, 1.89, 2.95, 4.21 Please explain your answer.

OpenStudy (vishal_kothari):

sin2x - cosx = 1 or, 2*sin x* cos x - cos x=1 or, cos x (2*sin x -1) =1 So, we have either cos x =1 or (2*sin x-1)= 1 If cos x =1 , then x = 0 degree If (2*sin x-1) =1, then 2*sin x = 1+1 or, sin x = 1 or, sin x = sin 90 So, x = 90 degree = 3.14/2 radian = 1.57